OpenStudy (anonymous):

Multiply && Simplify: (x+8)^2 a. x^2+16 b. x^2+16x+64 c. x^2+64 d. x^2+16+16

OpenStudy (anonymous):

C right ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(x+8)^2 = (x+8)(x+8)

jimthompson5910 (jim_thompson5910):

Hopefully that may help

jimthompson5910 (jim_thompson5910):

If not, then (x+8)(x+8) = x(x+8)+8(x+8) So (x+8)^2 = x(x+8)+8(x+8)

OpenStudy (anonymous):

@jim_thompson5910 : B

jimthompson5910 (jim_thompson5910):

yes you got it